The order reads 5% dextrose in0.45% sodium chloride to infuse Ml/ hr. The iv tubing reads 10drops / ml. What rate in drops/ min will the Rn regulate the iv fluid
Answer #1

Total volume×drop factor/Time in minutes.

Total volume=500ml

Time=60 min

Drop factor=10drops/ml

500×10/60=5000/60=83.33gtt/min.